# CLAP: Isolating Content from Style through Contrastive Learning with Augmented Prompts

### v5

@@ 4.1 Isolating Content from Style with Augmented Images
To generate training image data, we combine class names with various image and object attributes to create text prompts for each class. Using a stable diffusion model [39], we produce synthetic images that comprise our training dataset $\mathcal{D}_{x}$. The creation of template prompts for stable diffusion is based on attributes such as object size, color, image type, and art style. As detailed in table 1, the attributes include 10 colors and 3 object sizes, as well as 8 image types and 2 art styles. By assembling these attributes into prompts such as "a [art style] [image type] of a [object size] [object color] [class]", we generate 480 unique texts per class, from which one image per prompt is synthesized. Further details on image synthesis and examples are provided in § B.1.
@@
Table 1: Template-based prompts. Attributes used to generate text prompts follow the structured format: "a [art style] [image type] of a [object size] [object color] [class]", where "[class]" represents the class names.
@@
yellow, green, black, blue, multicolored, orange, red, white, brown, purple large, small, normal sized painting, cartoon, infograph, sketch, photograph, clipart, mosaic art, sculpture realistic, impressionistic
@@
For the image augmentation procedures, we adopt techniques commonly used in contrastive learning [7, 41, 8], specifically random cropping and color distortion.
Disentangled Network Structure
@@
Since the training process is based on CLIP’s pre-trained lower-dimensional features, our disentangled network adopts a multi-layer perceptron (MLP) architecture. To fully benefit from the pre-trained CLIP text encoder, we construct a residual MLP featuring a zero-initialized projection, acting as the disentangled network, as depicted in fig. 3. This design enables learning directly from the pre-trained representation space, avoiding a random starting point. It is inspired by ControlNet’s zero-conv operation [46], which we adapt to a zero-linear operation within our residual MLP.
Within this architecture, the main branch includes a zero-initialized, bias-free linear layer, positioned after a SiLU activation and a normally initialized linear layer. Conventionally, the dimensions of features before the initial linear layer, between the first and second linear layers, and after the second linear layer are referred to as the input ($d_{\text{in}}$), latent ($d_{\text{mid}}$), and output ($d_{\text{out}}$) dimensions, respectively.
@@
To rectify any mismatches between the input and output dimensions, the network employs nearest-neighbor downsampling in the shortcut path, thereby ensuring both dimensional alignment and the preservation of sharpness in the input features. During inference, a weighting parameter $\alpha>0$ is introduced to modulate the contribution of features from the main branch before integration with the input features, whereas this parameter remains fixed at 1 during training.
Inference
@@
After training, the disentangled network $\mathbf{f}^{*}_{c}$ is utilized following CLIP’s image encoder to extract visual content features. Moreover, given that vision-language data generation is rooted in a unified latent space, as described in § 3, $\mathbf{f}^{*}_{c}$ can be seamlessly integrated with CLIP’s image and text encoders to enhance zero-shot capabilities. As shown in fig. 2(c), for an image $\mathbf{x}$, the operation is formulated as the composition function $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{x}(\mathbf{x})$, and similarly, for a text $\mathbf{t}$, as $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t})$. This integration preserves CLIP’s zero-shot functionality while providing refined features through improved disentanglement of content.
@@
Figure 3: Structure of the disentangled network. The architecture encompasses a residual block featuring a zero-initialized, bias-free linear layer to initiate optimization from the input feature space. When the input and output dimensions differ, a downsampling operation is employed to ensure alignment. During inference, a scalar parameter $\alpha$ balances the contributions of the main branch and the input features prior to their combination.
@@ 4.2 Isolating Content from Style with Augmented Prompts
Despite progress in disentangling content and style via image augmentation, adequately altering all style factors in an image remains challenging due to the high dimensionality and complexity of style information in images. Achieving substantial style changes through augmentation—essential for complete disentanglement [41]—is difficult with existing image augmentation techniques. On the contrary, text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. To further explore the disentanglement of content, we propose Contrastive Learning with Augmented Prompts (CLAP).
@@
As depicted in fig. 2(b), CLAP employs an InfoNCE loss to train a disentangled network atop CLIP’s pre-trained text encoder, keeping the encoder’s gradients fixed—similar to Im.Aug. Leveraging the simpler structure of text, the template-based prompts previously utilized for synthesizing images now serve as the training text dataset, denoted by $\mathcal{D}_{t}$. Utilizing the same disentangled network as in Im.Aug, the learning objective of CLAP is defined as:
@@
$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{t}_{i}\}_{i=1}^{b}\in\mathcal{D}_{t}}\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},\tau)+\lambda\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}^{c}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},1),$ (4)
@@
where $\mathbf{f}^{*}_{t}$ denotes the pre-trained CLIP text encoder. The term $\mathbf{t}_{i}$ refers to a text prompt from $\mathcal{D}_{t}$, and $\mathbf{\tilde{t}}_{i}$ represents its augmented view, produced via prompt augmentation techniques. On the right-hand side of the equation, $\mathbf{t}_{i}^{c}$ denotes the class name associated with the text prompt $\mathbf{t}_{i}$. This strategy aims to enhance variation between prompt pairs, especially in cases where the text dataset $\mathcal{D}_{t}$ contains a limited number of samples. The hyperparameter $\lambda$ controls the relative weight of the second term in the overall loss function. All other symbols in Eq. 4 retain their definitions as described earlier.
@@
After training, the learned disentangled network is seamlessly integrated with both of CLIP’s encoders to extract content representations, as shown in fig. 2(c).
Prompt Augmentation
To ensure text prompts undergo stylistic changes without compromising their content, we develop specific augmentation techniques for synthetic text prompts. Drawing inspiration from Easy Data Augmentation (EDA) [42], we adapt the Random Deletion (RD) and Random Swap (RS) techniques, customizing them for our structured prompt format. To avoid inadvertently altering the content—such as by introducing new object names or modifying the core semantics—we exclude random insertions and replacements.
@@
Our primary augmentation techniques include Object Size Deletion (OSD), Object Color Deletion (OCD), Image Type Deletion (ITD), Art Style Deletion (ASD), and Swapping Prompt Order (SPO). Each is applied with a fixed probability, as detailed in table 2. Additionally, for downstream datasets with few categories, we enrich the training set population using an additional augmentation method named IGN (Inserting Gaussian Noise). Following initialization protocols from prompt learning [48, 47], we insert zero-mean Gaussian noise with a standard deviation of 0.02 and noise length of 4 into the tokenized prompts.
@@
Intuitively, these prompt augmentation methods parallel random masking techniques used in image augmentation [17, 6]. However, prompt augmentations are often more effective and precise than their image-based counterparts. This is because prompt augmentations can directly target and remove specific style elements without affecting core content, whereas image masking operates at the pixel or patch level and may inadvertently distort content or fail to sufficiently alter style.
@@
Table 2: Prompt augmentation techniques. Various augmented views are generated from an original text prompt using specific techniques: OSD (Object Size Deletion), OCD (Object Color Deletion), ITD (Image Type Deletion), ASD (Art Style Deletion), and SPO (Swapping Prompt Order).
Prompt Variant OSD OCD ITD ASD SPO
a realistic painting of a large red car a realistic painting of a red car a realistic painting of a large car a realistic of a large red car a painting of a large red car a large red car in a realistic painting
@@ 5 Experiments
We conduct three primary experiments to assess our method:
  1. 1.
    Zero-shot evaluation with diverse prompts to gauge performance and robustness under prompt perturbations;
  2. 2.
    Linear probe tests on few-shot samples to evaluate the efficacy of the learned representations in few-shot settings;
  3. 3.
    Adversarial attack assessments on zero-shot and one-shot classifiers to determine their resistance to adversarial threats.
We further conduct an ablation study on hyperparameters, explore the impact of different prompt augmentation combinations and various sources of training prompts on CLAP’s performance, and replicate experiments across different CLIP model sizes.
@@ 5.1 Experimental Setup
Implementation
Im.Aug and CLAP are implemented using the ViT-B/16 CLIP model and executed on an NVIDIA RTX 3090 GPU. To ensure reproducibility, the random seed for all stochastic processes is fixed at 2024. Additional implementation details are provided in § A.1.
Datasets
CLAP is evaluated on four multi-domain datasets to assess its performance across diverse environments: PACS [23] (4 domains, 7 categories), VLCS [1] (4 domains, 5 categories), OfficeHome [37] (4 domains, 65 categories), and DomainNet [35] (6 domains, 345 categories). For conciseness, we present average results across domains for each dataset. Detailed results for individual domains are provided in § A.4.
Compute Efficiency
CLAP demonstrates faster convergence and shorter training times compared to Im.Aug. Specifically, CLAP completes training on the PACS and VLCS datasets in approximately 11 minutes, OfficeHome in about 14 minutes, and DomainNet in roughly 47 minutes. In contrast, Im.Aug requires around 16 minutes for PACS and VLCS, 50 minutes for OfficeHome, and 3.3 hours for DomainNet. Both Im.Aug and CLAP retain CLIP’s inference efficiency due to the disentangled network’s compact two-layer MLP structure.
@@ 5.2 Main Results
Zero-Shot Performance
To assess zero-shot capabilities, CLAP is evaluated using three specific fixed prompts: ZS(C), which uses only the class name as "[class]"; ZS(PC), formatted as "a photo of a [class]"; and ZS(CP), structured as "a [class] in a photo". To further examine robustness, we introduce a dynamic prompt, ZS(NC), structured as "[noise][class]", where "[noise]" refers to Gaussian noise with mean 0 and standard deviation 0.02.
As shown in table 3, CLAP outperforms both CLIP and Im.Aug across all prompts and datasets. Unlike the uniform improvement achieved by CLAP, Im.Aug yields inconsistent results. Notably, CLAP exhibits a substantial advantage over CLIP when evaluated with the dynamic ZS(NC) prompt, demonstrating its effectiveness in improving zero-shot performance through enhanced content disentanglement.
Table 3: Zero-shot performance across prompt variants. Fixed prompts: "C" for [class], "CP" for a [class] in a photo, "PC" for a photo of a [class]; dynamic prompt "NC" is [noise][class] with Gaussian noise. CLAP consistently outperforms CLIP, while image augmentation (Im.Aug) shows mixed results.
Prompt Method        PACS      VLCS OfficeHome DomainNet Overall
@@
To evaluate robustness to prompt perturbations, we analyze the variance in zero-shot performance across prompts using the range ($R$) and standard deviation ($\delta$) from ZS(C), ZS(CP), and ZS(PC). We also measure the performance drop ($\Delta_{(NC)}$) from ZS(C) to ZS(NC) to assess resilience to noisy prompts.
As shown in table 4, CLAP significantly reduces zero-shot variance, exhibiting lower $\delta$ and $R$ values, and a smaller drop in performance under noise, compared to both Im.Aug and CLIP. While Im.Aug offers modest variance reduction, it is clearly less effective than CLAP. These results demonstrate CLAP’s superior robustness in maintaining consistent performance across prompt variations.
Table 4: Prompt robustness analysis. CLAP more effectively reduces zero-shot performance variance across prompts than image augmentation. $R$ and $\delta$ denote the range and standard deviation across ZS(C), ZS(CP), and ZS(PC); $\Delta_{(NC)}$ indicates the performance drop from ZS(C) to ZS(NC). Lower values imply greater robustness.
Metric Method        PACS      VLCS OfficeHome DomainNet Overall
@@
Few-Shot Performance
We conduct evaluations using 1-shot, 4-shot, 8-shot, and 16-shot linear probes across each domain within the four datasets. As illustrated in fig. 4, CLAP significantly outperforms both CLIP and Im.Aug in few-shot learning scenarios. Notably, in the 1-shot setting, CLAP achieves performance improvements over the linear-probe CLIP model by margins of +10%, +3.5%, +2.5%, and +1.5% on the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively. These improvements are especially pronounced when compared to the gains observed with Im.Aug, highlighting CLAP’s superior effectiveness in few-shot learning. Detailed quantitative results are provided in § A.4.
Figure 4: Few-shot linear probe comparisons of image-encoder features show that CLAP enhances CLIP’s few-shot performance more effectively than Im.Aug. "ZS" indicates zero-shot performance using a [class] prompt.
Adversarial Performance
To evaluate adversarial robustness, we test both zero-shot (ZS(C)) and one-shot classifiers against three popular adversarial attack methods: FGSM [15], PGD [30], and CW [5]. Adversarial samples are generated with 1 iteration for FGSM and 20 iterations for both PGD and CW, using an $\epsilon$ of 0.031.
As shown in table 5, classifiers leveraging CLAP representations exhibit significantly greater robustness than those based on CLIP. Across the four datasets, CLAP’s zero-shot and one-shot classifiers outperform CLIP by +7.6% and +8.5% against FGSM, +1.0% and +11.7% against PGD-20, and +1.1% and +2.3% against CW-20, respectively. These gains surpass the improvements achieved by Im.Aug: +4.4% and +4.6% (FGSM), +0.3% and +6.2% (PGD-20), and 0% and +1.3% (CW-20). These results indicate that CLAP enhances robustness against adversarial attacks in both zero-shot and one-shot scenarios.
Table 5: Adversarial robustness evaluation. Both Im.Aug and CLAP improve the robustness of CLIP’s zero-shot ([class]) and 1-shot classifiers under adversarial attacks (FGSM, PGD-20, CW-20). CLAP consistently achieves the highest resilience across all datasets.
Setting Method FGSM PGD-20 CW-20 Avg.

### v6

@@ 4.1.1 Data Synthesis and Image Augmentation
To generate training image data, we combine class names with various image and object attributes to create text prompts for each class. Using a stable diffusion model [39], we produce synthetic images that comprise our training dataset $\mathcal{D}\mathbf{x}$. The creation of template prompts for stable diffusion is based on attributes such as object size, color, image type, and art style. As detailed in Tab. 1, the attributes include 10 colors and 3 sizes for objects, and 8 types and 2 art styles for images. By assembling these attributes into prompts like "a [art style] [image type] of a [object size] [object color] [class]", we generate 480 unique texts for each class, from which one image per prompt is synthesized. Further details on image synthesis and examples are available in Sec. 0.B.1.
@@
Table 1: Template-based prompts. Attributes used to generate text prompts follow the structured format "a [art style] [image type] of a [object size] [object color] [class]", where "[class]" represents the class names.
@@
yellow, green, black, large,  painting, cartoon, infograph, realistic,
blue, multicolored, orange, small, sketch, photograph, clipart,  impressionistic
red, white, brown, purple  normal sized mosaic art, sculpture
@@
For the image augmentation procedures, we adopt techniques commonly used in contrastive learning practice [7, 41, 8], specifically random cropping and color distortion.
4.1.2 Disentangled Network Structure
@@
Since the training process is based on CLIP’s pre-trained lower-dimensional features, our disentangled network adopts a multi-layer perceptron (MLP) architecture. To fully benefit from the pre-trained CLIP text encoder, we construct a residual MLP featuring a zero-initialized projection, acting as the disentangled network, as depicted in Fig. 3. This design enables learning directly from the pre-trained representation space, avoiding a random starting point, inspired by ControlNet’s zero-conv operation [46], which we adapt to a zero-linear operation within our residual MLP.
@@
Within this architecture, the main branch includes a zero-initialized, bias-free linear layer positioned subsequent to the combination of a SiLU activation and a normally initialized linear layer. Conventionally, the dimensions of features before the initial linear layer, situated between the first and second linear layers, and following the second linear layer, are named as the input $d_{in}$, latent $d_{mid}$, and output $d_{out}$ dimensions, respectively. To rectify any mismatches between the input and output dimensions, the network employs nearest-neighbor downsampling within the shortcut path, thereby ensuring both alignment and the preservation of sharpness for the input features. During the inference stage, a weighting parameter $\alpha>0$ is introduced to modulate the portion of features emanating from the main branch before their integration with the input features, whereas this parameter remains constant at 1 throughout the training phase.
4.1.3 Inference
@@
After training, the disentangled network $\mathbf{f^{*}_{c}}$ is utilized following CLIP’s image encoder to extract visual content features. Moreover, given that vision-language data generation is rooted in a unified latent space, as depicted in Sec. 3, $\mathbf{f^{*}_{c}}$ can be seamlessly integrated with CLIP’s image and text encoders to enhance zero-shot capabilities. As shown in Fig. 2(c), for an image $\mathbf{x}$, the operation is formulated as the composition function $\mathbf{f^{*}_{c}\circ f^{*}_{x}(x)}$, and similarly, for a text $\mathbf{t}$, as $\mathbf{f^{*}_{c}\circ f^{*}_{t}(t)}$. This integration preserves CLIP’s zero-shot functionality while achieving refined features through the improved disentanglement of content.
@@
Figure 3: Structure of the disentangled network. The architecture encompass a residual block featuring a zero-initialized, bias-free linear layer to commence optimization from the input feature space. When the input and output dimension differ, a downsampling operation is utilized to achieve alignment. During inference, a scalar parameter $\alpha$ balance the main branch and input features before combination.
@@ 4.2 Isolating Content from Style with Augmented Prompts
Despite progress in disentangling content and style via image augmentation, adequately altering all style factors in an image remains challenging due to the high dimensionality and complexity of style information in images. Achieving substantial style changes through augmentation, essential for complete disentanglement [41], is difficult with existing image augmentation techniques. On the contrary, text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. To further exploring the disentanglement of content, we propose Contrastive Learning with Augmented Prompts (CLAP).
@@
As depicted in Fig. 2(b), CLAP employs an InfoNCE loss to train a disentangled network atop CLIP’s pre-trained text encoder, keeping the encoder’s gradients fixed, similar to Im.Aug. Leveraging the simpler structure of text, the template-based prompts previously utilized for synthesizing images now serve as the training text dataset, denoted by $\mathcal{D}_{\mathbf{t}}$. Utilizing the same disentangled network as in Im.Aug, the learning objective of CLAP is outlined as follows:
@@
$\mathbf{f^{*}_{c}}=\mathop{\mathrm{argmin}}_{\mathbf{f_{c}}}\mathop{\mathbb{E}}_{\{\mathbf{t}_{i}\}_{i=1}^{b}\in\mathcal{D}_{\mathbf{t}}}\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{t}};\{\mathbf{t}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},\tau)+\lambda\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{t}};\{\mathbf{t}^{c}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},1),$ (4)
@@
where $\mathbf{f}^{*}_{\mathbf{t}}$ denotes the pre-trained CLIP text encoder. The term $\mathbf{t}_{i}$ references a text prompt from $\mathcal{D}\mathbf{t}$, and $\mathbf{\tilde{t}}_{i}$ represents its augmented view, produced via prompt augmentation techniques. On the equation’s right side, $\mathbf{t}_{i}^{c}$ specifies the class name associated with the text prompt $\mathbf{t}_{i}$. This strategy aims to enhance variations between prompt pairs, especially in cases where the text dataset $\mathcal{D}_{\mathbf{t}}$ has a very limited number of samples. Here, $\lambda$ serves for adjusting the second term’s importance in the total loss function. All other symbols in Eq. 4 maintain their definitions as described earlier.
@@
After training, the learned disentangled network is seamlessly integrated with both of CLIP’s encoders to extract content representations, as depicted in Fig. 2(c).
4.2.1 Prompt Augmentation
@@
To ensure text prompts undergo stylistic changes without compromising their content, we have developed specific augmentation techniques for synthetic text prompts. Drawing inspiration from Easy Data Augmentation (EDA) techniques [42], we adapted the Random Deletion (RD) and Random Swap (RS) techniques from EDA, customizing them to suit our prompt structure. To avoid inadvertently altering the content by introducing new object names or changing the core idea of a text prompt, our augmentation methods do not include random word insertions or replacements. Our primary augmentation techniques are Object Size Deletion (OSD), Object Color Deletion (OCD), Image Type Deletion (ITD), Art Style Deletion (ASD), and Swapping Prompt Order (SPO), each applied with a certain probability, as detailed in Tab. 2. Additionally, for down-stream datasets with few categories, to rich the population of training samples, we use an additional augmentation, named IGN (Inserting Gaussian Noise). Following the initializing protocol of prompt learning methods [48, 47], we insert a zero-mean Gaussian noise with 0.02 standard deviation with a noise length equals to 4, to the tokenized prompts.
@@
Intuitively, these prompt augmentation methods parallel random masking techniques used in image augmentation [17, 6]. However, prompt augmentations are more effective and precise than their image counterparts. This effectiveness arises because prompt augmentations can specifically target and eliminate a particular style element without impacting the content, whereas image masking, operating at the pixel or patch level, might inadvertently damage content information or lead to insufficient style changes.
@@
Table 2: Prompt augmentation techniques. Various augmented views are generated from an original text prompt using specific augmentation techniques: OSD (Object Size Deletion), OCD (Object Color Deletion), ITD (Image Type Deletion), ASD (Art Style Deletion), and SPO (Swapping Prompt Order).
Original OSD OCD ITD ASD SPO
a realistic a realistic a realistic a realistic a painting a large red
painting of a  painting of a  painting of a of a of a car in a
large red car red car large car  large red car  large red car  realistic painting
@@ 5 Experiments
We conduct three primary experiments to assess our method: (1) zero-shot evaluation with diverse prompts to gauge zero-shot performance and its robustness to prompt perturbations; (2) linear probe tests on few-shot samples to evaluate the efficacy of the learned representations in few-shot settings; and (3) adversarial attack assessments on zero-shot and one-shot classifiers to determine their resistance to adversarial threats. We further conduct an ablative study on hyper-parameters, explore the impact of different prompt augmentation combinations and various sources of training prompts on CLAP’s performance, and replicate experiments across different CLIP model sizes.
@@ 5.1 Experimental Setup
Implementation. Im.Aug and CLAP are implemented using the ViT-B/16 CLIP model and executed on an NVIDIA RTX 3090 GPU. To ensure reproducibility, the random seed for all stochastic processes is fixed at 2024. More information on implementation details is provided in Sec. 0.A.1.
Datasets. CLAP is assessed across four multi-domain datasets to examine its performance in varied environments: PACS [23] (4 domains, 7 categories), VLCS [1] (4 domains, 5 categories), OfficeHome [37] (4 domains, 65 categories), and DomainNet [35] (6 domains, 345 categories). For conciseness, we present average results across the domains for each dataset. Detailed experimental outcomes for each domain within these datasets are provided in Sec. 0.A.4.
Compute efficiency. CLAP demonstrates faster convergence and shorter training times compared to Im.Aug. For CLAP, training on the PACS and VLCS datasets is completed in roughly 11 minutes, OfficeHome in approximately 14 minutes, and DomainNet in about 47 minutes. In contrast, Im.Aug requires around 16 minutes for PACS and VLCS, 50 minutes for OfficeHome, and 3.3 hours for DomainNet. Both Im.Aug and CLAP maintain CLIP’s inference efficiency due to the disentangled network’s efficient two-layer MLP structure.
@@ 5.2 Main Results
5.2.1 Zero-Shot Performance
To assess zero-shot capabilities, CLAP undergoes evaluation using three specific fixed prompts: ZS(C), utilizing only the class name within "[class]"; ZS(PC), with the format "a photo of a [class]"; and ZS(CP), structured as "a [class] in a photo". To thoroughly examine zero-shot proficiency, a dynamic prompt, ZS(NC), formatted as "[noise][class]", is also used, where "[noise]" signifies the introduction of Gaussian noise characterized by a mean of 0 and a standard deviation of 0.02.
As presented in Tab. 3, CLAP surpasses both CLIP and Im.Aug across all evaluated prompts for every dataset. Unlike the uniform enhancement in zero-shot performance CLAP achieves over CLIP, Im.Aug displays inconsistent results. A closer examination reveals CLAP’s superiority over CLIP is especially significant for the dynamic ZS(NC) prompt. This demonstrates CLAP’s effectiveness in significantly improving zero-shot performance compared to the original CLIP representations.
Table 3: Zero-shot results across three distinct prompts: "C" for "[class]", "CP" for "a [class] in a photo", "PC" for "a photo of a [class]", and a dynamic prompt "NC" for "[noise][class]" showcase that CLAP consistently outperforms CLIP’s zero-shot performance across all datasets, whereas image augmentation exhibits mixed outcomes.
Prompt Method Zero-shot performance, avg. top-1 acc. (%) (↑)
 PACS  VLCS Off.Home  Dom.Net Overall
@@ 5.2.1 Zero-Shot Performance
In assessing the model’s robustness to prompt perturbations, we examine the variances in zero-shot performance across different prompts by analyzing the range ($R$) and standard deviation ($\delta$) of results derived from ZS(C), ZS(CP), and ZS(PC). Additionally, we investigate the decline ($\Delta_{(NC)}$) in performance from ZS(C) to ZS(NC) as a broad indicator of resilience to noised prompts.
As presented in Tab. 4, CLAP significantly reduces the variance in zero-shot performance across various testing prompts, evidenced by markedly lower values of $\delta$ and $R$, and a less pronounced decrease in performance with a noised prompt, in contrast to Im.Aug and the baseline representations of CLIP. Although Im.Aug aids in reducing performance variance to some extent, its efficacy is notably inferior to that of CLAP. These findings highlight CLAP’s enhanced robustness in maintaining consistent zero-shot performance across a diverse array of prompts.
Table 4: CLAP more effectively reduces zero-shot performance variance across prompts than image augmentation, with $R$ and $\delta$ indicating the range and standard deviation for ZS(C), ZS(CP), and ZS(PC). The decrease $\Delta_{(NC)}$ from ZS(C) to ZS(NC) highlights CLAP’s enhanced robustness against prompt perturbations.
Metric Method Performance variance, avg. top-1 acc. (%) (↓)
 PACS  VLCS Off.Home  Dom.Net Overall
@@
5.2.2 Few-Shot Performance
We conduct evaluations of 1-shot, 4-shot, 8-shot, and 16-shot linear probes across each domain within the four datasets. As illustrated in Fig. 4, CLAP significantly outperforms both CLIP and Im.Aug in few-shot learning scenarios. Notably, in the 1-shot setting CLAP achieves performance improvements over the linear-probe CLIP model by margins of +10%, +3.5%, +2.5%, and +1.5% on the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively. These improvements are especially significant in comparison to the gains observed with Im.Aug counterparts, underpinning CLAP’s efficacy in few-shot scenarios. For detailed quantitative results, please refer to Sec. 0.A.4.2.
Figure 4: Few-shot linear probe comparisons of image-encoder features show that CLAP enhances CLIP’s few-shot performance more effectively than Im.Aug. In the accompanying figure, "ZS" indicates the zero-shot performance using a "[class]" prompt.
5.2.3 Adversarial Performance
To assess adversarial robustness, zero-shot (ZS(C)) and one-shot classifiers are evaluated against prominent adversarial attack methods, such as FGSM [15], PGD [30], and CW [5], by generating adversarial samples for testing. For FGSM, 1 adversarial iteration is employed, whereas for PGD and CW, 20 iterations are used, all with an epsilon of 0.031. As indicated in Tab. 5, classifiers utilizing CLAP representations demonstrate superior resilience to these adversarial attacks compared to those based on CLIP representations. Across the four datasets, CLAP’s zero-shot and 1-shot classifiers surpass CLIP by margins of +7.6% and +8.5% against FGSM, +1.0% and +11.7% against PGD-20, and +1.1% and +2.3% against CW-20, respectively. These figures notably exceed the performance improvements of +4.4% and +4.6% against FGSM, +0.3% and +6.2% against PGD-20, and 0% and +1.3% against CW-20 achieved by Im.Aug. The result suggests that CLAP efficiently enhances robustness against adversarial attacks in both zero-shot and one-shot scenarios.
Table 5: Image augmentation and CLAP both enhance CLIP’s zero-shot with the "[class]" prompt and 1-shot robustness against adversarial attacks, with CLAP showing greater improvements.
Setting Method Avg. top-1 acc. (%) under adversarial attacks(↑)
FGSM PGD-20 CW-20  Avg.